OpenStudy (anonymous):

Verify tan^2(x/2) = secx-1/secx+1

OpenStudy (tkhunny):

\(\tan^{2}(x/2) = \sec(x) - \dfrac{1}{\sec(x)} + 1\) Are you SURE that's what you meant? That IS what you wrote.

OpenStudy (anonymous):

\[\tan ^{2}(\frac{ x }{2 }) = \frac{ secx-1 }{secx+1 }\]

OpenStudy (tkhunny):

I see. Please remember your Order of Operations. Use parentheses to clarify. Hmmm... \(\cos(2x) = 2\cos^{2}(x) - 1 = 1 - 2\sin^{2}(x)\) Transforming... \(\cos(x) = 2\cos^{2}(x/2) - 1 = 1 - 2\sin^{2}(x/2)\) Solving for the squared half angles: \(\cos^{2}(x/2) = \dfrac{\cos(x)+1}{2}\) \(\sin^{2}(x/2) = \dfrac{\cos(x)-1}{-2} = \dfrac{1 - \cos(x)}{2}\) Did that get us anywhere?

OpenStudy (anonymous):

Not really I just completely don't understand this stuff at all have been trying for 3 hours and am about ready to give up

OpenStudy (tkhunny):

You're breaking my heart, here. Do you know this one: \(\tan(x) = \dfrac{\sin(x)}{\cos(x)}\)?